# Adapters for Altering LLM Vocabularies: What Languages Benefit the Most?

### v1

@@ 5.3 Evaluation
Figure 2: Average scores of original Mistral and its adaptation with new vocabulary, only replacing embeddings and fixing the body of transformer modules. “-multi” indicates models with a multilingual vocabulary, which includes five languages covering all languages with two separate models, while “-mono” refers to models with monolingual vocabulary. Our VocADT adaptation outperforms the original Mistral and matches or exceeds ZeTT. xx-en and en-xx indicate MT tasks.
@@ 6.1 Overall Task Performance
We first present the controlled comparison on diverse tasks of the original Mistral with new vocabulary variants obtained by our vocabulary adaptation approach (VocADT) and the ZeTT and OFA baselines. Figure 2 presents the average performance across multiple multilingual MT, NLI, reasoning, and QA tasks. Language-wise results are in Appendix B. Overall, adapting the vocabulary using VocADT generally leads to better performance compared to the original Mistral model, and either surpasses or performs on par with ZeTT. MMMLU is the only task where Mistral still holds the top spot; however, the performance gap between the new and original embeddings is smaller with VocADT approach than with ZeTT. Remarkably, VocADT-multi achieves these results with only two models for the eight languages tested, whereas ZeTT requires a separate model for each language.
6.2 Which Languages Benefit the Most from Vocabulary Adaptation?
Instead of “averaging out” the languages, we get closer look on the impact of vocabulary adaptation in a language-wise manner. Figure 4 shows the increased rate of task performance and tokenization statistics for various languages after applying different vocabulary adaptation methods. The results are shown for Latin group languages (Swahili, Indonesian, Estonian, and Haitian Creole) and Mixed group (Korean, Greek, Russian, and Bulgarian). We compare VocADT and ZeTT against the original Mistral model. The vertical axis of the increase rate is fixed to the positive range to see the benefit trends more easily and all the numbers including the negative range of the increase rate are in Appendix B. We use the FLORES development set for counting the tokens by various tokenizers where the semantic contents of every language are the same.
In Figures 4(a), we observe that Latin script languages consistently benefit from vocabulary adaptation, regardless of the task, adaptation method employed. However, even non-Latin languages show improvements when they suffer from severe over-fragmentation, as seen in the case of Greek in Figures 4(b). Among the eight languages, Greek is the most fragmented by the Mistral tokenizer, and it demonstrates significant improvement after the adaptation to less fragmented vocabulary, particularly in MT tasks, while other non-Latin languages in Mixed group show zero or modest gains.
@@ 6.2 Which Languages Benefit the Most from Vocabulary Adaptation?
Figure 4: Effect of vocabulary adaption on mitigating over-fragmentation and task performance. The $y$-axis for the increase rate on the left side is limited to the positive range. Languages with Latin scripts or those experiencing severe fragmentation benefit the most. xx-en and en-xx are machine translation tasks.
@@ 6.3 Does Script Matter for Language grouping?
Figure 5 compares the performance and token count reduction between two non-English grouping strategies: Mixed-script (ko, el, ru, bg) and Cyrillic-script (ru, bg, uk, kk) languages. The results are shown for the common languages of Russian and Bulgarian across several tasks. For Russian, the consistent script language group performs better, especially in the MT task, though the improvement rate is not substantial. For Bulgarian, both grouping strategies deliver nearly identical results. In summary, the results suggest that maintaining a consistent script within a group enhances performance, though overall outcomes seem to be influenced more by the language type itself than by the grouping strategy.
6.4 Is the Auxiliary Loss Helpful?
We examine the effects of the auxiliary loss that aims to mitigate the divergence of original status for overlapping words as described in Section 3.3. Figure 6 illustrates the impact on task performance of vocabulary adaptation with and without auxiliary loss on Latin and Mixed group vocabulary. We report one language in each group along with English.
For the Mixed group plus English vocabulary (right plot of Figure 6), maintaining the embedding values of overlapping words shows slight effectiveness. However, for Latin languages (left plot of Figure 6), applying auxiliary loss ($\alpha=0.1$) leads to a significant decline in overall task performance. We hypothesize that non-Latin languages are less prone to have word collisions with the original vocabulary compared to the Latin group, as the Mistral model is largely English (Latin) centric. As a result, retaining the established embeddings for overlapped words in Latin group vocabulary and Mistral vocabulary may disrupt effective learning due to the possible similarity in scripts with English. On the other hand, keeping the original embeddings during the adaptation for overlapping tokens in Mixed may be helpful to maintain the already established embeddings for overlapped tokens while adjusting the embeddings for new non-Latin script tokens.
Figure 6: Effects of the auxiliary loss on various settings. Keeping the embedding of overlapping words close to the original is ineffective for Latin groups but is effective in Mixed vocabulary.
@@ 6.4 Is the Auxiliary Loss Helpful?
7 Impact of Vocabulary Adaptation on Downstream Fine-tuning
@@ 8 Conclusion
We propose a simple and effective vocabulary adaptation method using a vocabulary adapter that addresses key challenges in existing approaches. Our approach consistently outperforms the original Mistral model after the adaptation phase across various tasks and after the following full-weight finetuning on machine translation. Furthermore, our methods prove more effective than, or at least on par with, strong vocabulary adaptation baselines, without relying on external embeddings or language constraints, offering a flexible and scalable solution for handling multiple languages. Our experiments cover a wide range of languages and scripts, revealing that languages with Latin scripts or severe fragmentation benefit the most, and that maintaining the embedding of overlapping words proves helpful for non-Latin languages but is less effective for Latin groups. We also explored different grouping strategies, finding that while maintaining consistent scripts within a group offers some benefits, the impact is relatively minor. Lastly, with a focus on machine translation, we confirm that vocabulary adaptation remains effective even after fine-tuning, with VocADT emerging as the most effective approach.
Appendix A Details of Training
@@
Appendix B Language-wise Results of Vocabulary Adaptations
@@
Table 9: Tables of various vocabulary adaptation methods.

### v2

@@ 6.1 Overall Task Performance
We first present the controlled comparison on diverse tasks of the original Mistral with new vocabulary variants obtained by our vocabulary adaptation approach (VocADT) and the ZeTT and OFA baselines. Figure 2 presents the average performance across multiple multilingual MT, NLI, reasoning, and QA tasks. Language-wise results are in Appendix E. Overall, adapting the vocabulary using VocADT generally leads to better performance compared to the original Mistral model, and either surpasses or performs on par with ZeTT. MMMLU is the only task where Mistral still holds the top spot; however, the performance gap between the new and original embeddings is smaller with VocADT approach than with ZeTT. Remarkably, VocADT-multi achieves these results with only two models for the eight languages tested, whereas ZeTT requires a separate model for each language.
@@
Figure 4: Effect of vocabulary adaption on mitigating over-fragmentation and task performance. The $y$-axis for the increase rate on the left side is limited to the positive range. Languages with Latin scripts or those experiencing severe fragmentation benefit the most. xx-en and en-xx are machine translation tasks. See Appendix E for individual task performance values.
6.2 Which Languages Benefit the Most from Vocabulary Adaptation?
Instead of “averaging out” the languages, we get closer look on the impact of vocabulary adaptation in a language-wise manner. Figure 4 shows the increased rate of task performance and tokenization statistics for various languages after applying different vocabulary adaptation methods. The results are shown for Latin group languages (Swahili, Indonesian, Estonian, and Haitian Creole) and Mixed group (Korean, Greek, Russian, and Bulgarian). We compare VocADT and ZeTT against the original Mistral model. The vertical axis of the increase rate is fixed to the positive range to see the benefit trends more easily and all the numbers including the negative range of the increase rate are in Appendix E. We use the FLORES development set for counting the tokens by various tokenizers where the semantic contents of every language are the same.
In Figures 4(a), we observe that Latin script languages consistently benefit from vocabulary adaptation, regardless of the task, adaptation method employed. However, even non-Latin languages show improvements when they suffer from severe over-fragmentation, as seen in the case of Greek in Figures 4(b). Among the eight languages, Greek is the most fragmented by the Mistral tokenizer, and it demonstrates significant improvement after the adaptation to less fragmented vocabulary, particularly in MT tasks, while other non-Latin languages in Mixed group show zero, modest, or even negative gains.
In Appendix B, we further discuss the pronounced performance declines observed in Korean compared to Russian or Bulgarian within the same Mixed group. Despite improving fragmentation for Korean, we suspect that the linear combination assumption is insuffiencnt given the lack of representation of Korean characters in the original vocabulary.
@@ 6.3 Does Script Matter for Language grouping?
Figure 5 compares the performance and token count reduction between two non-English grouping strategies for Russian and Bulgaian: Mixed-script (ko, el, ru, bg) and Cyrillic-script (ru, bg, uk, kk) languages. For Russian, the consistent script language group performs slightly better, especially in the MT task. For Bulgarian, both grouping strategies deliver nearly identical results. Overall, the results suggest that maintaining a consistent script within a group enhances performance, though outcomes tend to beinfluenced more by the language type itself than by the grouping strategy.
6.4 Scalability and Generalizability of VocADT
We further explore the language scalability of the method with All language groups including 11 languages (§F.1), and the generalizability of our VocADT findings to other language models (§F.2). Both scalability and generalizability experiments show that the All group follows trends similar to the Latin, Mixed, and Cyrillic setups while the performance trends observed with LLaMA are consistent with those seen in Mistral.
7 Impact of Vocabulary Adaptation on Downstream Fine-tuning
@@ 8 Conclusion
We propose a simple and effective vocabulary adaptation method using a vocabulary adapter. Our approach consistently outperforms the original Mistral model after the adaptation phase across various tasks and after the following full-weight finetuning on machine translation. Furthermore, our method is on par with or more effective than strong vocabulary adaptation baselines, without relying on external embeddings or language constraints, offering a flexible and scalable solution for handling multiple languages. Our experiments cover a wide range of languages and scripts, revealing that languages with Latin scripts or severe fragmentation benefit the most. We also explored different grouping strategies, finding that maintaining consistent scripts within a group offers relatively minor benefits. Lastly, with a focus on machine translation, we confirm that vocabulary adaptation remains effective even after full weight fine-tuning, and VocADT is the most effective approach.
Acknowledgements
We thank Anthony Aue for early discussions, and Marcin Junczys-Dowmunt and the anonymous reviewers for their insightful and helpful feedback.
Appendix A Is the Auxiliary Loss Helpful?
We examine the effects of the auxiliary loss that aims to mitigate the divergence of original status for overlapping words as described in Section 3.3. Figure 6 illustrates the impact on task performance of vocabulary adaptation with and without auxiliary loss on Latin and Mixed group vocabulary. We report the average of four non-English languages in each group along with English.
For Latin languages (left plot of Figure 6), omitting the auxiliary loss ($\alpha=0$) performs slightly better or comparably to using a non-zero $\alpha$. For the Mixed group plus English vocabulary (right plot of Figure 6), maintaining the embedding values of overlapping words shows slight effectiveness in both non-English and English. We hypothesize that non-Latin languages are less prone to have word collisions with the original vocabulary compared to the Latin group, as the Mistral model is largely English (Latin) centric. As a result, retaining the established embeddings for overlapped words in Latin group vocabulary and Mistral vocabulary may disrupt effective learning due to the possible similarity in scripts with English. On the other hand, keeping the original embeddings during the adaptation for overlapping tokens in Mixed may be helpful to maintain the already established embeddings for overlapped tokens while adjusting the embeddings for new non-Latin script tokens.
Figure 6: Effects of the auxiliary loss on various settings.
This auxiliary loss can be thought of as a form of regularization. There is a long history of applying regularization during adaptation of MT models to control the adaptation process by limiting the amount that the output distribution or the weights of the fine-tuned model can vary from the original model weights. Prior work has explored using dropout and L2 regularization (Miceli Barone et al., 2017), cross entropy (Khayrallah et al., 2018), freezing parts of the network (Wuebker et al., 2018; Thompson et al., 2018), and Elastic Weight Consolidation (Kirkpatrick et al., 2017; Thompson et al., 2019a; b). As our auxiliary loss has mixed effectiveness depending on language characteristics, future work could consider other methods.
Appendix B Discussions on non-alphabetic scripts and Possible Limitations of Linear Combination Assumption
As shown in Figure 4(b), the performance decline for Korean in VocADT vocabulary transfer (-14% to -15% in MT) is more pronounced than for Russian or Bulgarian (-1% to -4%) within the same Mixed group, even though the fragmentation improvement for Korean is greater. This deviates from the expected performance improvements seen in Greek, where mitigating extreme over-fragmentation (150k tokens) led to gains. Although Korean has significant over-fragmentation (70k tokens), its severity is less than half that of Greek and closer to Russian and Bulgarian (54k tokens).
One possible explanation is the limitation of our assumption that new embeddings can be solely represented as linear combinations of old embeddings. This assumption may not hold well for Korean, which uses the non-alphabetic Hangul script. In Hangul, tokens represent entire syllables or consonant-vowel combinations rather than individual phonemes, making them difficult to decompose into subwords using the original tokenizer. For instance, the new token “처럼” (meaning “like” in English) appears frequently in Korean. However, the original Mistral vocabulary lacks a dedicated token for “럼”, preventing proper decomposition of “처럼” without resorting to byte-level tokens. These byte-level fallbacks may not effectively capture the linguistic structure of the character, potentially degrading performance. This issue is less prevalent in alphabetic scripts such as Latin, Cyrillic, or Greek, where words can be easily broken down into individual characters. This limitation may account for the observed performance discrepancy.
Appendix C Details of Training, Baseline & Evaluation
@@
For parallel training data for the MT task, we use bitext from the NLLB dataset (Schwenk et al., 2021b; Heffernan et al., 2022; NLLB Team et al., 2022) [^10] https://huggingface.co/datasets/allenai/nllb This includes web-scraped data, which has the potential to include nosise such as text being automatically identified as the wrong language, mis-aligned or mis-translated segments, and low-quality machine translated segments (Khayrallah & Koehn, 2018; Caswell et al., 2020; Dodge et al., 2021; Kreutzer et al., 2022; Thompson et al., 2024a). We use LASER3 (Artetxe & Schwenk, 2019) to select higher quality segments for fine-tuning. LASER has been used extensively to both locate parallel segments on the web (Schwenk et al., 2021a; b) as well as for filtering noisy sentence and document pairs (Chaudhary et al., 2019; Koehn et al., 2020; Thompson & Koehn, 2020; Sloto et al., 2023).
In adapter training for VocADT, we use a (peak) learning rate of 2e-6 with a cosine scheduler, a maximum sequence length of 512 tokens, a warm-up ratio of 0.01, and a weight decay of 0.01. In full-weight fine-tuning phase, we mostly follow the training setting from ALMA.
Details of Baseline.
For ZeTT, we use multilingual hypernetwork for Mistral-7B. [^11] https://github.com/bminixhofer/zett We use the code of OFA [^12] https://github.com/cisnlp/ofa and FOCUS [^13] https://github.com/konstantinjdobler/focus to create new embeddings for Mistral-7B.
Machine Translation Metrics.
We assess translation quality using xCOMET-XL (Guerreiro et al., 2023), as recent WMT metric shared tasks (Freitag et al., 2023; 2024) have found neural metrics like Yisi (Lo, 2019; Lo & Larkin, 2020), Bert-score (Zhang et al., 2019), Prism (Thompson & Post, 2020a; b), Comet (Rei et al., 2020), BLEURT (Sellam et al., 2020) correlate much better with human judgements, than surface-form metrics like BLEU (Papineni et al., 2002) or chrF (Popović, 2015; 2017) which consider only surface form. Trained metrics like the comet and BLEURT, which train on prior human annotations of translation quality, achieve the highest correlation with human judgments. While these correlations are less strong out of domain (relative to the domains used in WMT, e.g. FLORES) the trained metrics still outperform surface level ones (Zouhar et al., 2024).
We also caveat that xCOMET-XL does not consider context when judging translation quality, and context has been shown to be an important aspect of translation quality evaluation (for a thorough overview see (Castilho & Knowles, 2024), especially for LLMs Karpinska & Iyyer (2023). While there have been several efforts to incorporate context in MT evaluation (e.g. Vernikos et al. (2022); Deutsch et al. (2023); Raunak et al. (2024), there is no consensus in the community as to which method, so we stick to the established xCOMET-XL at the sentence level. Finally, metric differences, especially small ones, may not correspond to statistically significant differences (Koehn, 2004; Deutsch et al., 2021; Lo et al., 2023; Thompson et al., 2024b).
Appendix D Computational Cost VocADT
We report the computational cost of our approach. We use a batch size of 128 (four A100 GPUs * 8 batch size * 4 gradient accumulation) and a sequence length of 512. The FLOPs per token for VocADT is 17.7GFLOPs/token, resulting in 1160T “FLOPs per batch” (128 * 512 * 17.7G). Our training requires 38k update steps (2.49B, roughly 0.5B per langs). Therefore, the computational cost of a VocADT model (1160T “FLOPs per batch” x 38k step). We use profile() method of accelerator Python library and our estimation of the FLOPs per token for Mistral-7B is 14.2 GFLOPs/token.
Appendix E Language-wise Results of Vocabulary Adaptations
@@
Table 9: Tables of various vocabulary adaptation methods. The works in bold linearly combine original embeddings to generate new embeddings.
@@
Appendix F Additional Experiment for Scalability and Generalizability of VocADT
F.1 Combining Languages of Latin, Mixed, and Cyrillic into All group
In Section 6.3, we observed that while grouping languages for the new vocabulary with a consistent script improves performance, script-based grouping strategies had little overall impact. This suggests that we can enhance the method’s scalability for greater practicality with minimal performance tradeoffs. In this section, we explore a multilingual group with shared vocabulary at larger scales to provide better insights into scalability for multilingual setups.
We combine languages from the MultiLatin, Mixed, and Cyrillic —groups into one unified set into All. This set comprises 11 languages—English and 10 non-English languages (Swahili, Indonesian, Estonian, Haitian, Korean, Greek, Russian, Bulgarian, Ukrainian, and Kazakh) as listed in Table 1. Following our experimental setup of 0.5B tokens per language, we train on a combined corpus of 5.5B monolingual tokens, covering all 11 languages. [^14] Available in https://huggingface.co/h-j-han/Mistral-7B-VocADT-50k-All We set $\alpha=0$.
Tables 10 and 11 show that the All group follows trends similar to the initial Latin, Mixed, and Cyrillic setups. Figure 7 further illustrates that while the token count for the All group is slightly higher than that of the Multi group setup, it remains significantly lower than that of the original Mistral model.
F.2 Generalization to LlaMA
We primarily conducted our experiments using the Mistral model. To validate the generalizability of our VocADT findings to other language models, we also test our approach on an additional candidate LM, LLaMA (Touvron et al., 2023).
We conducted an additional adaptation experiment using LLaMA2-7B, following the same experimental setup described in the main section. Figure 7 shows that the severity of fragmentation in LLaMA is similar to that in Mistral, with Greek being the most severely fragmented language followed by Korean. Tables 10 and 11 confirm that the performance trends observed with LLaMA are consistent with those seen in Mistral. Overall, Latin group languages benefit largely from vocabulary adaptation, while non-Latin languages in the Mixed group show minus or modest gains, except for Greek, which benefits due to its severe fragmentation. These findings validate that our method generalizes effectively to another language model.
Table 10: xx-en and en-xx MT results with xCOMET-XL score and the increase rate from the original Mistral after the vocabulary adaptation—only replacing embeddings while fixing the rest. The tables compare the All 11-language group versus the Multi groups—Latin, Mixed, and Cyrillic (each comprising 5 languages). We also compare the experiments using Mistral versus LLaMA as the base model.
MT xx-en Mistral Llama
Orig VocADT-multi VocADT-all Orig VocADT-multi
Total # of Models 1 3 1 1 3
sw-en 0.485 0.801 65.32% 0.775 59.89% 0.359 0.698 94.43%
id-en 0.946 0.942 -0.44% 0.919 -2.89% 0.954 0.933 -2.20%
et-en 0.722 0.899 24.46% 0.851 17.79% 0.496 0.858 72.98%
ht-en 0.554 0.669 20.72% 0.63 13.74% 0.392 0.645 64.54%
ko-en 0.882 0.755 -14.39% 0.834 -5.41% 0.872 0.776 -11.01%
el-en 0.438 0.760 73.59% 0.856 95.44% 0.439 0.777 76.99%
ru-en 0.959 0.927 -3.33% 0.929 -3.14% 0.951 0.93 -2.21%
bg-en 0.952 0.918 -3.56% 0.92 -3.38% 0.941 0.916 -2.66%
Avg (8 pairs) 0.742 0.834 12.35% 0.839 13.03% 0.675 0.817 21.04%
uk-en 0.944 0.915 -3.07% 0.909 -3.74% 0.947 0.897 -5.28%
kk-en 0.411 0.763 85.82% 0.751 82.92% 0.286 0.611 113.64%
Avg (10 pairs) 0.729 0.835 14.49% 0.837 14.76% 0.664 0.804 21.08%
MT en-xx Mistral Llama
Orig VocADT-multi VocADT-all Orig VocADT-multi
en-sw 0.238 0.562 135.88% 0.466 95.48% 0.291 0.367 26.12%
en-id 0.778 0.837 7.65% 0.763 -1.89% 0.868 0.872 0.46%
en-et 0.309 0.643 108.37% 0.587 90.12% 0.279 0.581 108.24%
en-ht 0.308 0.329 7.03% 0.312 1.40% 0.286 0.315 10.14%
en-ko 0.703 0.598 -14.99% 0.631 -10.23% 0.669 0.566 -15.40%
en-el 0.384 0.413 7.56% 0.635 65.56% 0.297 0.511 72.05%
en-ru 0.900 0.854 -5.17% 0.855 -5.02% 0.877 0.824 -6.04%
en-bg 0.899 0.859 -4.43% 0.854 -4.96% 0.826 0.825 -0.12%
Avg (8 pairs) 0.565 0.637 12.77% 0.638 12.98% 0.549 0.608 10.75%
en-uk 0.865 0.851 -1.59% 0.83 -4.05% 0.83 0.814 -1.93%
en-kk 0.222 0.522 135.11% 0.555 150.05% 0.188 0.354 88.30%
Avg (10 pairs) 0.560 0.647 15.40% 0.649 15.79% 0.541 0.603 11.46%
Table 11: XNLI, XCOPA, Belebele, and MMMLU results with Accuracy score and the increase rate from the original Mistral after the vocabulary adaptation—only replacing embeddings while fixing the rest. The tables compare the All 11-language group versus the Multi groups—Latin, Mixed, and Cyrillic (each comprising 5 languages). We also compare the experiments using Mistral versus LLaMA as the base model.
XNLI Mistral Llama
Orig VocADT-multi VocADT-all Orig VocADT-multi
Total # of Models 1 3 1 1 3
en 0.550 0.553 0.47% 0.53 -3.64% 0.554 0.568 2.53%
sw 0.353 0.398 12.63% 0.397 12.46% 0.348 0.378 8.62%
el 0.419 0.387 -7.60% 0.396 -5.49% 0.370 0.382 3.24%
ru 0.488 0.490 0.48% 0.494 1.23% 0.425 0.47 10.59%
bg 0.425 0.457 7.63% 0.435 2.35% 0.424 0.388 -8.49%
Avg (5 langs) 0.447 0.457 2.24% 0.451 0.89% 0.424 0.437 3.00%
XCOPA Mistral Llama
Orig VocADT-multi VocADT-all Orig VocADT-multi
sw 0.510 0.574 12.55% 0.54 5.88% 0.522 0.546 4.60%
id 0.584 0.608 4.11% 0.592 1.37% 0.628 0.604 -3.82%
et 0.470 0.538 14.47% 0.5 6.38% 0.488 0.538 10.25%
ht 0.514 0.548 6.61% 0.538 4.67% 0.506 0.526 3.95%
Avg (4 langs) 0.520 0.567 9.14% 0.542 4.33% 0.536 0.5535 3.26%
Belebele Mistral Llama
Orig VocADT-multi VocADT-all Orig VocADT-multi
en 0.843 0.833 -1.18% 0.824 -2.29% 0.482 0.456 -5.39%
sw 0.391 0.440 12.50% 0.454 16.08% 0.262 0.289 10.31%
id 0.647 0.638 -1.38% 0.636 -1.65% 0.380 0.346 -8.95%
et 0.439 0.538 22.53% 0.54 23.03% 0.312 0.319 2.24%
ht 0.397 0.507 27.72% 0.522 31.59% 0.287 0.322 12.20%
ko 0.666 0.616 -7.52% 0.644 -3.25% 0.336 0.354 5.36%
el 0.442 0.566 27.90% 0.631 42.70% 0.301 0.357 18.60%
ru 0.727 0.696 -4.29% 0.71 -2.30% 0.428 0.378 -11.68%
bg 0.674 0.698 3.47% 0.694 2.91% 0.398 0.392 -1.51%
Avg (9 langs) 0.581 0.614 5.83% 0.629 8.33% 0.354 0.357 0.86%
uk 0.728 0.693 -4.76% 0.682 -6.32% 0.398 0.352 -11.50%
kk 0.364 0.442 21.36% 0.427 17.18% 0.261 0.277 6.00%
Avg(11 langs) 0.574 0.606 5.50% 0.615 7.08% 0.349 0.349 0.05%
MMMLU Mistral Llama
Orig VocADT-multi VocADT-all Orig VocADT-multi
en 0.607 0.577 -4.88% 0.561 -7.53% 0.452 0.415 -8.19%
id 0.468 0.410 -12.49% 0.444 -5.17% 0.367 0.296 -19.35%
ru 0.500 0.468 -6.39% 0.468 -6.33% 0.355 0.34 -4.23%
Avg (3 langs) 0.525 0.485 -7.62% 0.491 -6.45% 0.391 0.351 -10.23%
uk 0.489 0.462 -5.57% 0.463 -5.34% 0.346 0.328 -5.20%
Avg (4 langs) 0.516 0.479 -7.14% 0.484 -6.18% 0.38 0.345 -9.21%
Figure 7: Token count reduction with new vocabulary. Each bar displays two percentage reduction values: the first (e.g., -53.30% in Swahili) indicates the reduction relative to the original Mistral model, while the second (e.g., -52.88%) represents the reduction relative to the original LLaMA model. We use the FLORES development set for counting the tokens by various tokenizers where the semantic contents of every language are the same. While All group with all 11 languages is slightly higher than that of the Multi group with five languages, it remains significantly lower than that of the original models. The severity of fragmentation in LLaMA is similar to that in Mistral, with Greek being the most severely fragmented language followed by Korean.